OpenStudy (anonymous):

Media services charges $20 for a phone and 30$ a month for its economy plan.The equation c=30t + 20 describes the total cost of c of operating media services phone for t months.The total cost of 10 months service is?

jimthompson5910 (jim_thompson5910):

c=30t + 20 c=30(10) + 20 c=300+20 c=320 So the total cost for 10 months of service is $320

OpenStudy (anonymous):

and I understand it better the way you broke it down,thanks again.